OpenStudy (anonymous):

How do you find the base when you only know the exponent? Is there a rule I don't know? Say you have something like this: r^8 = 6561 How would you find r?

OpenStudy (johnweldon1993):

You would take the 8th root of both sides...to isolate r r^8 = 6561 \[\sqrt[8]{r}^8\] = \[\sqrt[8]{6561}\]see how that eliminates the ^8 power?

OpenStudy (jkristia):

take log on both sides and bring down the exponent 8 log(r) = log (6561) log (r) = log (6561) / 8 r = 10^(log(6561) / 8)

OpenStudy (jkristia):

ah, of course @johnweldon1993 approach is easier

OpenStudy (johnweldon1993):

Yes but BOTH ways are a correct way of doing this :)

OpenStudy (anonymous):

Didn't understand much but maybe I'll get it later. Thanks!

OpenStudy (johnweldon1993):

well, what didn't you understand?

OpenStudy (anonymous):

I didn't really understand how you got to \[8\sqrt{6561}\] and I'm also not sure how to find the actual number from that.

OpenStudy (jkristia):

\[r^2 = 9\] now take the square root on both sides \[\sqrt{r^2} = \sqrt{9}\] square root and square cancels, so you have \[r = \sqrt{9}\]

OpenStudy (anonymous):

Hmm but why \[r^{2}\] and not \[r ^{8}\]

OpenStudy (jkristia):

it was just an example, and because square root is easier when giving an example (like sqrt(9) = 3). But it is the same principal.

OpenStudy (anonymous):

Yes, but \[\sqrt{r ^{8}}\] wouldn't be r, right? :/ that's the bit that confuses me.

OpenStudy (jkristia):

Correct, \[\sqrt{r^8} \neq r, but \sqrt [8]{r^8} = r\] In my example I used r^2 not r^8 just because it is easier. But if you have something your square e.g. (r^2) = 100, then to find r you will need to do the inversion which is taking the root, in this case (just an example) \[r = \sqrt{100} = 10\] I hope I did not confuse you

OpenStudy (anonymous):

So does that apply to all radical problems of that form? I mean -> \[\sqrt{a ^{b}} = \sqrt[b]{a ^{b}} = a\] all the time? I forgot lots of basic rules...

OpenStudy (jkristia):

no that is not correct, squarertoot is not the same as the b'th root. \[3^2 = 9 => \sqrt{9} = 3\] \[3^3 = 27 => \sqrt[3]{27} = 3\] but \[\sqrt{27} = 5.19615...\]

OpenStudy (anonymous):

This is one of those things that will probably take me ages to get ><

OpenStudy (jkristia):

I'm sure you will get it with a bit of practice. Maybe just remember 10 SQUAREed = 100, the SQUAREroot of 100 is 10. Nah, that probably doesn't make it any easier